# Memory Decoder: A Pretrained, Plug-and-Play Memory for Large Language Models

### v1

@@ 5.3 Cross-Model Adaptation
Table 4 demonstrates Memory Decoder’s exceptional plug-and-play capabilities across diverse model sizes and architectures. A single Memory Decoder (0.5B parameters) consistently enhances performance across all models in both the Qwen2 and Qwen2.5 families, spanning from 0.5B to 72B parameters. For smaller models like Qwen2-0.5B, our approach achieves dramatic perplexity reductions—transforming domain-specific performance from near-baseline levels to state-of-the-art results on both biomedical and financial text. Even for the largest models in the family, Memory Decoder provides substantial improvements, demonstrating that retrieval-augmented knowledge remains valuable regardless of model scale. Most impressively, a 0.5B model augmented with our Memory Decoder surpasses vanilla 72B models across all evaluated domains, achieving over 140× parameter efficiency. These results validate Memory Decoder’s core strength: a single pretrained memory component can enhance multiple models sharing the same tokenizer, providing efficient domain adaptation that scales from the smallest to the largest models while consistently outperforming existing approaches.
@@ 5.4 Cross-Vocabulary Adaptation
5.5 Knowledge-Intensive Reasoning Tasks
NQ HotpotQA
Llama3-8B 23.64 25.14
+ kNN-LM 24.00 (+0.36) 24.48 (-0.66)
+ MemDec (1B) 28.01 (+4.37) 27.72 (+2.58)
Table 5: Performance on knowledge-intensive QA tasks. Memory Decoder provides substantial improvements while kNN-LM shows minimal gains or degradation.
While retrieval-augmented methods excel at improving factual recall, they often struggle with tasks requiring both knowledge retrieval and complex reasoning. Prior work (Geng et al., 2024) has shown that kNN-LM can actually harm performance on knowledge-intensive QA tasks, despite retrieving from relevant Wikipedia corpora.
To evaluate Memory Decoder’s ability to handle such challenging scenarios, following the experimental setup of Geng et al. (2024), we trained a 1B parameter Memory Decoder on the same large heterogeneous corpus (see detailed corpus composition in Appendix C) and evaluated on Natural Questions (NQ) and HotpotQA benchmarks.
As shown in Table 5, Memory Decoder achieves substantial improvements on both benchmarks, in stark contrast to kNN-LM which shows marginal improvement on NQ and degradation on HotpotQA. Our approach successfully enhances the model’s ability to access factual knowledge without compromising its reasoning capabilities—addressing a fundamental limitation of traditional retrieval methods. These results demonstrate that by learning to internalize retrieval patterns rather than relying on explicit retrieval at inference time, Memory Decoder maintains the compositional reasoning abilities necessary for complex multi-hop questions while still benefiting from expanded knowledge access.
@@ 6.1 Case Study: Bridging Parametric and Non-Parametric Methods
Long-tail Knowledge: For factual information like "Jacobi" and "1906", Memory Decoder assigns dramatically higher probabilities than the base model (68.94% vs. 0.12% and 98.65% vs. 1.57%), successfully capturing the memorization benefits of non-parametric methods while far exceeding even kNN-LM’s retrieval capabilities.
@@
These observations confirm that Memory Decoder occupies a unique position: it enhances memorization of domain-specific and long-tail knowledge like non-parametric methods, while maintaining the generalization and reasoning capabilities inherent to parametric models. This best-of-both-worlds behavior explains Memory Decoder’s strong performance across diverse evaluation settings, from domain adaptation to knowledge-intensive reasoning tasks.
6.2 Sensitivity Analysis of Interpolation Parameter
$\alpha$ 0.40 0.45 0.50 0.55 0.60 0.65 0.70 0.75 0.80
Avg PPL 4.21 4.17 4.14 4.119 4.110 4.112 4.126 4.155 4.201
Relative 1.024 1.015 1.007 1.002 1.000 1.001 1.004 1.011 1.022
Table 7: Sensitivity analysis of interpolation parameter $\alpha$ across 11 Qwen models on the law domain. Performance remains stable within 2.5% across the entire range (0.4–0.8), demonstrating robustness to hyperparameter selection.
To assess the sensitivity of our approach to hyperparameter selection, we conducted comprehensive experiments varying the interpolation parameter $\alpha$ across all 11 Qwen models on the law domain. As shown in Table 7, Memory Decoder demonstrates remarkable robustness: performance varies by less than 2.5% across the entire $\alpha$ range (0.4–0.8), with optimal performance at $\alpha$=0.6. The relative performance row shows that even at the extremes of the range, perplexity increases by only 2.2% and 2.4% respectively. This stability is crucial for practical deployment, as it indicates that Memory Decoder does not require extensive hyperparameter tuning to achieve strong performance. The shallow degradation around the optimum further suggests that practitioners can confidently deploy Memory Decoder with minimal domain-specific calibration.
@@ 6.3 Impact of Memory Decoder Size
DAPT 15.47 12.78 11.10 10.16 12.38
+ MemDec-S (117M) 13.36 12.25 11.53 10.93 12.01
+ MemDec-M (345M) 12.08 11.59 10.92 10.43 11.26
+ MemDec-L (774M) 11.67 11.23 10.83 10.28 11.00
@@
Table 8 examines how Memory Decoder size affects performance across the GPT-2 family. Remarkably, even our smallest Memory Decoder (117M parameters) achieves competitive results with full-parameter DAPT—for instance, GPT2-medium with a small Memory Decoder outperforms GPT2-medium DAPT (12.25 vs. 12.78). As Memory Decoder size increases, performance consistently improves across all base models, with the large variant achieving the best average perplexity. These results validate that Memory Decoder provides an efficient alternative to full model fine-tuning: practitioners can choose the decoder size based on their computational constraints while maintaining the crucial advantage of preserving the original model’s capabilities.
6.4 Ablation on Pre-training Objective
Biomed
Qwen2.5-3B Qwen2.5-7B Qwen2-7B
+KL Only 3.93 3.84 3.86
+CE Only 4.81 4.66 4.69
+MemDec 3.64 3.57 3.59
Table 9: Ablation study on Memory Decoder’s pretraining objective components.
Table 9 compares our hybrid training approach against single-objective variants on the biomedical domain. Memory Decoder consistently outperforms both alternatives across all model configurations, with substantial improvements over both KL-only (distribution alignment) and CE-only (standard language modeling) training. Notably, the CE-only baseline corresponds to the DAPT setup, demonstrating that simply interpolating predictions with a domain-adapted model cannot achieve the same benefits as our integrated approach. More comprehensive DAPT model interpolation ablation can be found in Appendix E. The gap between our method and KL-only training highlights the importance of maintaining direct corpus supervision, while the improvement over CE-only validates the value of incorporating non-parametric distribution knowledge. These results confirm that combining both objectives enables Memory Decoder to capture diverse continuation patterns from retrieval-based distributions while preserving the structural patterns of the domain corpus.
@@ 7 Related Work
Retrieval-Augmented Generation (RAG) enhances language models by incorporating knowledge from external sources, with retrieval granularity ranging from documents (Chen et al., 2017) to passages (Guu et al., 2020; Lewis et al., 2020b; Izacard et al., 2023b) to tokens (Khandelwal et al., 2019b; He et al., 2021b; Min et al., 2022; Yogatama et al., 2021). Token-level retrieval achieves superior performance for rare patterns and out-of-domain scenarios but introduces substantial computation overhead during inference. While non-differentiable retrieval mechanisms prevent end-to-end optimization and memory token approaches (Chevalier et al., 2023) enable differentiable access but are limited to local contexts, Memory Decoder provides both differentiable optimization and full-dataset knowledge access without expensive retrieval operations or model-specific datastores.
@@
Domain adaptation techniques have evolved from domain-specific pre-training (SciBERT (Beltagy et al., 2019), BioBERT (Lee et al., 2020), ClinicalBERT (Huang et al., 2019)) to parameter-efficient methods like LoRA (Hu et al., 2022) and adapters (Wang et al., 2020; Diao et al., 2021, 2023). However, these approaches require model-specific modifications, preventing generalization across architectures. Memory Decoder addresses this limitation by providing a domain-specific memory module that enhances multiple frozen language models without parameter modifications, enabling cross-model adaptation within tokenizer families and efficient cross-tokenizer transfer with minimal additional training.
@@ 8 Conclusion
The key innovation of Memory Decoder lies in its versatility and efficiency. A single pretrained Memory Decoder can seamlessly enhance any model that shares the same tokenizer, and with minimal additional training, can be adapted to models with different tokenizers and architectures. This capability enables efficient domain adaptation across model families, dramatically reducing the resources typically required for specialized model development. Our results confirm that Memory Decoder preserves the performance benefits of retrieval-augmented methods while maintaining the general capabilities of the base models, avoiding the catastrophic forgetting often observed with parameter fine-tuning approaches.
@@ A.2 Downstream Performance
The general pattern centers around $\alpha$=0.3, which is consistent with the findings in Shi et al. (2022).
@@ A.3 Cross-Model and Cross-Vocabulary Adaptation
For domain-specific language modeling tasks (section 5.3 and 5.4), we tune $\alpha$ on the validation set by searching over $\{0.4,0.6,0.8,0.9\}$. We find that $\alpha=0.6$ yields the best performance in most cases.
@@ Appendix B Analysis of DAPT Performance on Downstream Tasks
Previous work has shown that domain-adaptive pretraining can adversely affect a model’s prompting ability (Cheng et al., 2023). Our experiments reveal that this effect is particularly pronounced when using domain-conditional PMI (DCPMI) scoring for evaluation, especially on tasks where label verbalizers overlap with the pretraining domain vocabulary.
@@
As shown in Table 12, while direct language modeling evaluation reveals only modest performance drops with DAPT, the DCPMI scores show dramatic degradation for smaller models. This discrepancy arises because we employ fuzzy verbalizers following Shi et al. (2022), and the label spaces for Yahoo and AGN tasks contain terms (e.g., “politics,” “technology”) that appear frequently in WikiText-103. When DAPT increases the domain probability for these terms, it causes the conditional PMI scores to drop substantially, as the denominator in the DCPMI calculation becomes inflated.
@@
Appendix C Knowledge-Intensive Reasoning Task Corpus Composition
For the knowledge-intensive reasoning experiments in Section 5.5, we constructed a large heterogeneous corpus following Geng et al. (2024). The corpus combines diverse text sources to provide broad coverage of factual knowledge and various writing styles. The complete corpus is publicly available at https://huggingface.co/datasets/wentingzhao/knn-prompt-datastore.
Corpus Size
WikiText-103 181MB
Amazon Reviews 89MB
$C$-NEWS 457MB
IMDB 45MB
Total 722MB
Table 13: Composition of the heterogeneous corpus used for training Memory Decoder on knowledge-intensive tasks.
Appendix D Domain-Specific Downstream Tasks
To comprehensively evaluate Memory Decoder’s ability to maintain in-context learning capabilities while achieving domain adaptation, we conducted extensive experiments on 13 real-world domain-specific tasks across biomedical, financial, and legal domains, following the same domain corpus and evaluation framework from Cheng et al. (2023). These benchmarks test both zero-shot and few-shot performance, providing a rigorous assessment of whether domain adaptation preserves the model’s prompting abilities—a known weakness of traditional DAPT approaches (Cheng et al., 2023).
D.1 Biomedical Domain Tasks
ChemProt MQP PubmedQA RCT USMLE Avg
(13-shot) (4-shot) (0-shot) (10-shot) (0-shot)
Qwen2.5-7B 24.40 83.44 63.70 70.10 36.92 55.71
+ DAPT 17.59 76.22 65.70 21.00 34.64 43.03
+ MemDec (0.5B) 24.40 84.09 64.40 74.06 36.84 56.76
Table 14: Performance on biomedical domain-specific tasks. Memory Decoder preserves in-context learning capabilities while DAPT shows severe degradation, particularly on RCT (70.10 → 21.00).
D.2 Financial Domain Tasks
FiQA_SA FPB Headline NER ConvFinQA Avg
(5-shot) (5-shot) (5-shot) (20-shot) (0-shot)
Qwen2.5-7B 80.46 70.96 87.08 68.92 60.53 73.59
+ DAPT 75.59 66.39 86.03 69.32 58.52 71.17
+ MemDec (0.5B) 81.34 71.25 87.95 69.21 63.69 74.68
Table 15: Performance on financial domain-specific tasks. Memory Decoder achieves consistent improvements across all tasks while DAPT degrades performance on most benchmarks.
D.3 Legal Domain Tasks
SCOTUS (micro) SCOTUS (macro) CaseHOLD (micro) CaseHOLD (macro) UNFAIR-ToS Avg
(0-shot) (0-shot) (0-shot) (0-shot) (4-shot)
Qwen2.5-7B 26.66 17.90 35.92 35.93 87.05 40.69
+ DAPT 28.33 16.82 35.70 35.69 87.18 40.74
+ MemDec (0.5B) 31.66 21.05 37.58 37.59 87.05 42.99
Table 16: Performance on legal domain-specific tasks. Memory Decoder shows substantial improvements on case-based reasoning tasks while maintaining strong performance on contract analysis.
D.4 Analysis
Across all three domains, Memory Decoder demonstrates a crucial advantage: it successfully enhances domain-specific performance while preserving the model’s in-context learning capabilities. DAPT, in contrast, shows significant degradation on many tasks, particularly those requiring few-shot or zero-shot reasoning. The most dramatic example occurs in the biomedical RCT task, where DAPT’s performance collapses from 70.10 to 21.00, while Memory Decoder improves it to 74.06.
This preservation of prompting abilities is essential for practical deployment, as real-world applications often require models to handle both domain-specific knowledge and general reasoning tasks. Memory Decoder achieves this balance by augmenting rather than modifying the base model, ensuring that the original capabilities remain intact while adding specialized domain knowledge through the learned retrieval patterns.
Appendix E Comparison with DAPT Model Interpolation
A natural baseline for domain adaptation is to train a small model with domain-adaptive pretraining (DAPT) and interpolate its predictions with the base LLM. This approach, while conceptually similar to Memory Decoder’s inference-time interpolation, relies on standard language modeling objectives rather than our novel distribution alignment training. To thoroughly evaluate this alternative, we conducted extensive comparisons between Memory Decoder and logit ensembling with DAPT models.
Base Model Baseline PPL + DAPT (Small) + MemDec (Small)
GPT2-small 24.89 15.95 13.36 (+2.59)
GPT2-medium 18.29 14.26 12.25 (+2.01)
GPT2-large 15.80 13.13 11.53 (+1.60)
GPT2-xl 14.39 12.30 10.93 (+1.37)
Average 18.34 13.91 12.01 (+1.90)
Table 17: Comparison between Memory Decoder and DAPT model interpolation on WikiText-103. Both approaches use the same small model architecture (124M parameters) but differ in training objectives.
As shown in Table 17, Memory Decoder consistently outperforms the DAPT model interpolation baseline across all GPT-2 variants, achieving an average improvement of 1.90 perplexity points. This gap is particularly pronounced for smaller base models, where Memory Decoder provides 2.59 perplexity improvement over DAPT interpolation for GPT2-small. These results validate that our hybrid training objective—combining distribution alignment with language modeling—enables superior memorization and retrieval of domain knowledge compared to standard DAPT approaches.
The performance difference stems from Memory Decoder’s ability to learn from non-parametric distribution patterns during training. While a DAPT model only learns to maximize likelihood on the domain corpus, Memory Decoder additionally learns to approximate the retrieval-based distributions from the datastore, capturing richer continuation patterns that better represent the domain knowledge. This finding aligns with our ablation results in Section 6.4, where the CE-only baseline (equivalent to DAPT) significantly underperforms our full method.
Furthermore, the consistent improvement across different model scales demonstrates that this advantage is not merely an artifact of model size or specific architectural choices, but rather a fundamental benefit of our training methodology. Even when applied to the largest GPT2-xl model, where DAPT already provides substantial gains, Memory Decoder achieves additional improvements of 1.37 perplexity points, confirming the complementary nature of distribution alignment and standard language modeling objectives.
Appendix F In-Context Learning Performance Analysis
Memory Decoder is designed as a complementary approach to existing adaptation methods, including in-context learning (ICL). While our primary comparisons focus on domain-adaptive pretraining due to our emphasis on learning from unlabeled corpora, it is crucial to verify that Memory Decoder preserves and potentially enhances the model’s ability to leverage in-context examples—a capability often compromised by traditional domain adaptation methods.
To thoroughly evaluate this aspect, we conducted experiments on the CaseHOLD legal reasoning benchmark with varying numbers of in-context examples.
0-shot 4-shot 8-shot 16-shot Best
CaseHOLD (micro)
Qwen2.5-7B 35.92 36.34 36.34 36.09 36.34
+ MemDec (0.5B) 37.58 (+1.66) 37.95 (+1.61) 38.29 (+1.95) 37.53 (+1.44) 38.29
CaseHOLD (macro)
Qwen2.5-7B 35.93 36.35 36.35 36.10 36.35
+ MemDec (0.5B) 37.59 (+1.66) 37.95 (+1.60) 38.30 (+1.95) 37.53 (+1.43) 38.30
Table 18: In-context learning performance on CaseHOLD legal reasoning task across different shot settings. Memory Decoder consistently improves performance, with peak performance at 8-shot.
Table 18 reveals several critical insights about Memory Decoder’s interaction with in-context learning:
Zero-shot superiority: Memory Decoder’s zero-shot performance (37.58/37.59) already surpasses the base model’s best few-shot performance (36.34/36.35), demonstrating that our method effectively encodes domain knowledge that would otherwise require multiple demonstration examples to achieve.
Complementary benefits: Memory Decoder continues to benefit from in-context examples, with performance improving from 0-shot (37.58) to 8-shot (38.29) settings. This improvement pattern confirms that our method preserves the model’s ability to learn from demonstrations while adding orthogonal domain knowledge.
Consistent advantages: Across all tested shot settings (0, 4, 8, 16), Memory Decoder maintains a stable improvement margin over the base model, ranging from 1.43 to 1.95 points. This consistency indicates that the domain knowledge encoded by Memory Decoder complements rather than interferes with in-context learning mechanisms.
Optimal shot selection: Both methods achieve peak performance at 8 shots, with slight degradation at 16 shots—likely due to noise from increased context length. Importantly, Memory Decoder’s pattern mirrors the base model’s, suggesting that our method does not introduce additional brittleness to long-context scenarios.
These results definitively establish that Memory Decoder not only provides strong zero-shot domain adaptation but also preserves and enhances the model’s ability to leverage in-context examples—a critical capability for practical deployment. This stands in stark contrast to DAPT, which, as shown in our domain-specific evaluations (Appendix D), often compromises in-context learning abilities while pursuing domain adaptation. The preservation of ICL capabilities makes Memory Decoder particularly valuable for real-world applications where models must handle both domain-specific tasks and general reasoning with varying amounts of available demonstrations.
Appendix G Characteristics of $k$-NN Distributions
G.1 Extreme Sparsity and Concentration

### v2

@@ 5.3 Cross-Model Adaptation
Table 4 demonstrates Memory Decoder’s exceptional plug-and-play capabilities across diverse model sizes and architectures. A single Memory Decoder (0.5B parameters) consistently enhances performance across all models in both the Qwen2 and Qwen2.5 families, spanning from 0.5B to 72B parameters. For smaller models like Qwen2-0.5B, our approach achieves dramatic perplexity reductions—transforming domain-specific performance to state-of-the-art results on both biomedical and financial text. Even for the largest models in the family, Memory Decoder provides substantial improvements, demonstrating that retrieval-augmented knowledge remains valuable regardless of model scale. These results validate Memory Decoder’s core strength: a single pretrained memory component can enhance multiple models sharing the same tokenizer, providing efficient domain adaptation that scales from the smallest to the largest models while consistently outperforming existing approaches.
@@ 6.1 Case Study: Bridging Parametric and Non-Parametric Methods
Long-tail Knowledge: For factual information like "Jacobi" and "1906", Memory Decoder assigns dramatically higher probabilities than the base model (68.94% vs. 0.12% and 98.65% vs. 1.57%), successfully capturing the memorization benefits of non-parametric methods.
@@
These observations confirm that Memory Decoder occupies a unique position: it enhances memorization of domain-specific and long-tail knowledge like non-parametric methods, while maintaining the generalization and reasoning capabilities inherent to parametric models.
@@ 6.2 Impact of Memory Decoder Size
DAPT 14.76 12.78 11.10 10.16 12.20
+ MemDec-small (124M) 13.36 12.25 11.53 10.93 12.01
+ MemDec-medium (345M) 12.08 11.59 10.92 10.43 11.26
+ MemDec-large (774M) 11.67 11.23 10.83 10.28 11.00
@@
Table 6 examines how Memory Decoder size affects performance across the GPT2 family. As Memory Decoder size increases, performance consistently improves across all base models, with the large variant achieving the best average perplexity. These results validate that Memory Decoder provides an efficient alternative to full model fine-tuning: practitioners can choose the decoder size based on their computational constraints while maintaining the crucial advantage of preserving the original model’s capabilities.
6.3 Ablation on the pre-training objective
Base Model Baseline PPL + DAPT-small + MemDec-small
GPT2-small 24.89 15.95 13.36 (-2.59)
GPT2-medium 18.29 14.26 12.25 (-2.01)
GPT2-large 15.80 13.13 11.53 (-1.60)
GPT2-xl 14.39 12.30 10.93 (-1.37)
Average 18.34 13.91 12.01 (-1.90)
Table 7: Memory Decoder vs. DAPT model interpolation on WikiText-103. Both use 124M parameter models with different training objectives.
We compare Memory Decoder against logit interpolation with a DAPT model. Table 7 shows Memory Decoder consistently outperforms DAPT interpolation across all GPT2 scales, with an average gain of 1.90 perplexity points. Notably, the gains persist even at larger scales, confirming that our hybrid pre-training objective provides complementary value beyond what standard language modeling objectives can achieve, regardless of model size.
@@ 7 Related Work
Retrieval-Augmented Generation (RAG) enhances language models by incorporating knowledge from external sources, with retrieval granularity ranging from documents [Chen et al., 2017] to passages [Guu et al., 2020, Lewis et al., 2020, Izacard et al., 2023] to tokens [Khandelwal et al., 2019, He et al., 2021, Min et al., 2022, Yogatama et al., 2021]. Token-level retrieval achieves superior performance for rare patterns and out-of-domain scenarios but introduces substantial computation overhead during inference. While non-differentiable retrieval mechanisms prevent end-to-end optimization and memory token approaches [Chevalier et al., 2023] enable differentiable access but are limited to local contexts, Memory Decoder provides both differentiable optimization and full-dataset knowledge access without expensive retrieval operations or model-specific datastores.
@@
Domain adaptation techniques have evolved from domain-specific pre-training (SciBERT [Beltagy et al., 2019], BioBERT [Lee et al., 2020], ClinicalBERT [Huang et al., 2019]) to parameter-efficient methods like LoRA [Hu et al., 2022] and adapters [Wang et al., 2020, Diao et al., 2021, 2023]. However, these approaches require model-specific modifications, preventing generalization across architectures. Memory Decoder addresses this limitation by providing a domain-specific memory module that enhances multiple frozen language models without parameter modifications, enabling cross-model adaptation within tokenizer families and efficient cross-tokenizer transfer with minimal additional training.
@@ 8 Conclusion
The key innovation of Memory Decoder lies in its versatility and efficiency. A single pretrained Memory Decoder can seamlessly enhance any model that shares the same tokenizer, and with minimal additional training, can be adapted to models with different tokenizers and architectures. This capability enables efficient domain adaptation across model families, dramatically reducing the resources typically required for specialized model development. Our results confirm that Memory Decoder preserves the performance benefits of retrieval-augmented methods while maintaining the general capabilities of the base models, avoiding the catastrophic forgetting often observed with fine-tuning approaches.
@@ 9 Limitations
Acknowledgement
This work is sponsored by the National Key Research and Development Program of China (No. 2023ZD0121402), Shanghai Fundamental Research Program for General AI Models (No. 2025SHZDZX0251101), and the National Natural Science Foundation of China (NSFC) grant (No.62576211).
@@ A.2 Downstream Performance
The general pattern centers around $\alpha$=0.3, which is consistent with the findings in Shi et al. [2022].
@@ A.3 Cross-Model and Cross-Vocabulary Adaptation
For domain-specific language modeling tasks (section 5.3 and 5.4), we tune $\alpha$ on the validation set by searching over $\{0.4,0.6,0.8,0.9\}$.
@@ Appendix B Analysis of DAPT Performance on Downstream Tasks
Previous work has shown that domain-adaptive pretraining can adversely affect a model’s prompting ability [Cheng et al., 2023]. Our experiments reveal that this effect is particularly pronounced when using domain-conditional PMI (DCPMI) scoring for evaluation, especially on tasks where label verbalizers overlap with the pretraining domain vocabulary.
@@
As shown in Table 10, while direct language modeling evaluation reveals only modest performance drops with DAPT, the DCPMI scores show dramatic degradation for smaller models. This discrepancy arises because we employ fuzzy verbalizers following Shi et al. [2022], and the label spaces for Yahoo and AGN tasks contain terms (e.g., “politics,” “technology”) that appear frequently in WikiText-103. When DAPT increases the domain probability for these terms, it causes the conditional PMI scores to drop substantially, as the denominator in the DCPMI calculation becomes inflated.
@@
Appendix C Characteristics of $k$-NN Distributions
C.1 Extreme Sparsity and Concentration